# Rustam Ibragimova,b    Jihyun Kimc    Anton Skrobotovd a Imperial College Business School b New Economic School

### v2

@@ 4.1 Predictive Regressions with Multiple Predictors
Consider model (7) and define $z_{t}=[\operatorname{sign}(x_{1,t}),\ldots,\operatorname{sign}(x_{K,t})]^{\prime}$. Suppose that for each $t\in\{1,\dots,T\}$, $\mathbb{E}[z_{t-1}u_{t}]=\mathbf{0}_{K\times 1}$, $0<\mathbb{E}[|x_{k,t-1}|]<\infty$ for all $k$, and that the matrix $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ is invertible. [^10] Even when $\operatorname{sign}(x_{t-1})$ is constant, the univariate Cauchy estimator and associated tests remain well defined. In this case, the estimator simplifies to $\check{\beta}=\sum_{t=1}^{T}y_{t}\big/\sum_{t=1}^{T}x_{t-1}$, implying $\check{\beta}-\beta=\sum_{t=1}^{T}u_{t}\big/\sum_{t=1}^{T}x_{t-1}$, and Proposition 3.4 continues to hold. Hence, the lack of sign variation does not affect the validity of our methods in the univariate case. In the multiple-predictor case, however, the invertibility of $\mathbb{E}[z_{t-1}x_{t-1}^{\prime}]$ imposes mild restrictions on sign changes across predictors. For practical applications, one can induce variation in the sign instrument by recentering the predictor, for example, $\tilde{x}_{t-1}=x_{t-1}-t^{-1}\sum_{s=1}^{t}x_{s-1}$, which preserves both the martingale structure and the asymptotic validity of the estimator. Then, the joint null hypothesis $H_{0}$ holds if and only if $H_{0}^{(k)}$ holds for all $k=1,\ldots,K$.
Lemma 4.1 justifies the use of multiple hypothesis testing based on univariate Cauchy estimators. [^11] See Harvey2015 for an application of the multiple-testing framework in predictive regressions, and KMS2015 for joint-predictability tests in the IVX framework. Note that the IVX approach may lose validity under heavy-tailed predictors or continuous-time data, whereas our method remains robust in such settings. In conjunction with the hybrid test introduced in Section 3.2, we compute the statistic $\tau(\check{\beta}_{k})$ for each parameter $\beta_{k}$, where $\check{\beta}_{k}$ denotes the corresponding Cauchy estimator. Let $p_{k}$ denote its $p$-value. The joint null hypothesis $H_{0}$ is rejected at level $\alpha$ if $\min_{k}p_{k}\leq\alpha/K$, following the Bonferroni correction.
This approach directly extends the univariate robust inference procedure to a multivariate setting and requires only mild conditions for the equivalence between $H_{0}$ and $\{H_{0}^{(k)}\}_{k=1}^{K}$. The Bonferroni correction imposes no assumptions on the joint distribution of the test statistics, which motivates its use here (see Holm1979, Holm1979; BenjaminiHochberg1995, BenjaminiHochberg1995; Shaffer1995, Shaffer1995).
We also note that if one additionally assumes that $\mathbb{E}[z_{t}z_{t}^{\prime}]$ is invertible, the joint hypothesis can be tested directly using a Wald-type statistic:
@@
and hence $W\to_{d}\chi_{K}^{2}$. [^12] As mentioned earlier, $z_{t}$ can be interpreted as an instrument. Therefore, one may use an alternative instrument, as in shephard2020, and construct a Wald-type test accordingly. We leave a systematic comparison between the Bonferroni-type multiple-testing procedure and the Wald-type joint test for future research.
@@ 4.2 Predictive Regressions with an Intercept
where $\bar{z}_{s}=s^{-1}\sum_{t=1}^{s}z_{t}$ for $z_{t}\in\{y_{t},x_{t-1},u_{t}\}$. However, $(u_{t}-\bar{u}_{T})$ is not an martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$, invalidating the martingale CLT used in Sections 2 and 3. Specifically, the Cauchy estimator becomes
@@
and estimate this first-differenced (FD) model on alternating subsets of observations. We focus on the even-indexed observations and define the modified Cauchy estimator:
@@
This estimator has two key properties. First, for even-indexed data, the regression error $u_{t}^{e}=u_{2t}-u_{2t-1}$ forms an MDS with respect to $\mathcal{F}_{t}^{e}:=\mathcal{F}_{2t}$ for $t=1,\ldots,T/2$. [^14] For odd-indexed data, $u_{t}^{o}=u_{2t+1}-u_{2t}$ forms an MDS with respect to $\mathcal{F}_{t}^{o}:=\mathcal{F}_{2t+1}$, yielding an analogous estimator $\check{\beta}_{o}$. Second, $\check{\beta}_{e}$ can again be viewed as an IV estimator, but it uses $\text{sign}(x_{t-2})$, which is $\mathcal{F}_{t-2}$-measurable, as the instrument. [^15] More generally, one may use $\text{sign}\!\left(\sum_{l\leq 2}w_{l}x_{t-l}\right)$ for deterministic weights $\{w_{l}\}$, provided $\mathbb{E}[\text{sign}(\sum_{l\leq 2}w_{l}x_{t-l})(x_{t-1}-x_{t-2})]\neq 0$. Hence, $\text{sign}(x_{2t-2})(u_{2t}-u_{2t-1})$ is an MDS with respect to $(\mathcal{F}_{t}^{e})$.
@@
Although the odd-indexed estimator $\check{\beta}_{o}$ has analogous properties, $\check{\beta}_{e}$ and $\check{\beta}_{o}$ are typically dependent, with the dependence structure determined by the DGP of $(x_{t})$. Hence, unless additional assumptions are imposed, we restrict attention to a single subset of observations—either with even or odd indices. [^16] Using only half of the data is not uncommon in predictive regressions. See, for example, ZhuCaiPeng2014 and Liu2019, who employ long-lag differencing to eliminate intercepts. In addition, Dufour uses a split-sample approach to address inference problems under a Markovian structure.
@@
$\mathbb{E}[\operatorname{sign}(x_{t-1})x_{t}]\neq\mathbb{E}[|x_{t-1}|].$
The condition may fail for certain unit-root processes. For instance, for a random walk $x_{t}=x_{t-1}+\varepsilon_{t}^{x}$, it does not hold. More generally, in the local-to-unity model of Phillips_Magdalinos_2007book,
@@
The first term diverges if and only if $\delta<1$ (see Lemma 3.2 of Phillips_Magdalinos_2007book, Phillips_Magdalinos_2007book). Since $\varepsilon_{t}^{x}$ and $x_{t-1}$ may be dependent, typically $\mathbb{E}[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0$, causing the second term to diverge as well. Hence, the condition fails only if $\delta=1$ and $\mathbb{E}[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]=0$. In all other cases ($0\leq\delta<1$ or nonzero covariance), the test remains consistent.
@@ 5 Finite Sample Performance
This section investigates the finite-sample performance of the proposed inference methods. Two sets of simulation experiments are conducted. The first set is based on a continuous-time model and compares our robust $t$-statistic–based tests, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and the hybrid test $\tau(\check{\beta})$ with the Cauchy RT test of CJP2016 and the Cauchy VC test of ibragimov2023new. The second set is based on a discrete-time predictive regression model with an intercept and compares our procedures with the IVX test of KMS2015.
@@ 5.1.1 Simulation Design
Following CJP2016 and ibragimov2023new, we consider the continuous-time predictive regression model
@@
where $V_{t}$ and $W_{t}$ are Brownian motions with $\mathbb{E}(V_{t}W_{t})=-0.98t$. The constant term in the predictive regression is set to zero without loss of generality, and recursive demeaning is applied. The model is observed at interval $\Delta=1/252$, corresponding to daily observations, so that a sample of length $T$ years contains $252T$ observations.
@@
  • GBM (Geometric Brownian motion): $d\sigma_{t}^{2}=\frac{1}{2}\frac{\bar{\omega}^{2}}{T}\sigma_{t}^{2}\,dt+\frac{\bar{\omega}^{2}}{\sqrt{T}}\sigma_{t}^{2}\,dZ_{t}$, where $Z_{t}$ is a Brownian motion correlated with $W_{t}$ such that $\mathbb{E}(W_{t}Z_{t})=-0.4t$ and $\bar{\omega}=9$.
We set $T\in\{5,20,50,100\}$ (corresponding to 60, 240, 600, and 1200 monthly observations) and $\bar{\kappa}\in\{0,5,10\}$ for the persistence parameter in (11), and consider a two-sided test of $H_{0}:\beta=0$ against $H_{A}:\beta\neq 0$.
@@ 5.1.2 Results
We first assess the empirical size of each test under the null hypothesis $\beta=0$. The results for the four volatility models (CNST, SB, RS, and GBM) and for $T\in\{5,20,50\}$ are reported in Table 1. Overall, both the $t$-statistic–based tests and the hybrid method exhibit satisfactory size performance, closely matching the nominal levels and performing comparably to the Cauchy RT and Cauchy VC tests. Among the $t$-based procedures, moderate partition numbers ($q=12$ or $q=16$) provide the most stable results, whereas smaller $q$ values tend to be mildly undersized. In the GBM case, where volatility is endogenously persistent, the $t$-statistic–based tests become slightly conservative but remain competitive with the Cauchy RT and VC methods.
Next, we analyze the finite-sample power properties of the tests. We consider $\beta\in\{0.004k,k=1,\cdots,5\}$ and $T\in\{20,50,100\}$ under the same volatility specifications. The results are summarized in Tables 25. The proposed tests exhibit power comparable to that of the Cauchy RT and Cauchy VC procedures. For small samples ($T=20$), the Cauchy RT and VC tests occasionally show higher power, but the difference diminishes as $T$ increases. In certain settings, our methods even outperform the existing approaches. For instance, $t_{16}(\check{\gamma})$ dominates under $\beta=0.02$, $\bar{\kappa}=0.5$, and regime-switching volatility (Table 4), whereas the hybrid test $\tau(\check{\beta})$ performs best under $\beta=0.004$, $\bar{\kappa}=20$, $T=20$, and regime-switching volatility.
@@ 5.2.1 Simulation Design
We now examine the finite-sample performance of the proposed tests in a discrete-time setting with an intercept, comparing them to the IVX test of KMS2015. The data-generating process (DGP) is specified as
$y_{t}$ $=\beta x_{t-1}+\sigma_{t}\varepsilon_{t},$
$x_{t}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{t-1}+\sigma_{t}\eta_{t},$ (12)
for $t=2,\ldots,T$, where $T\in\{60,240,600\}$ corresponds to 5, 20, and 50 years of monthly data. We set $\beta\in\{0.5k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider a one-sided test of $H_{0}\!:\beta=0$ against $H_{A}\!:\beta>0$. [^17] The IVX test of KMS2015 performs well in two-sided testing for a broad class of models. However, as shown in DEMETRESCU2023105271, the IVX method exhibits severe size distortions in one-sided tests when regressors are highly persistent and endogenous. For this reason, we focus on the one-sided case to demonstrate the performance of our methods in this setting.
@@
$\eta_{t}=\sum_{j=1}^{q}C_{j}v_{t-j},$
where $(v_{t},\eta_{t})$ are jointly normal with correlation 0.98-0.98. For the MA(2) case, $C_{1}=C_{2}=1/\sqrt{2}$ and $C_{q}=0$ for $q>3$; for the MA(4) case, $C_{1}=C_{2}=C_{3}=C_{4}=1/2$ and $C_{q}=0$ for $q>5$. The volatility process $\sigma_{t}$ follows the same specifications as in the continuous-time simulations, except that the GBM model is excluded.
We implement the hybrid tests based on the even and odd observations, denoted by $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, respectively (see (9)), along with their $t$-statistic versions (see (4)). For comparison, we also include the IVX test of KMS2015.

### v3

@@ 4.1 Predictive Regressions with Multiple Predictors
Consider model (S.1) and define $z_{t}=[\operatorname{sign}(x_{1,t}),\ldots,\operatorname{sign}(x_{K,t})]^{\prime}$. Suppose that for each $t\in\{1,\dots,T\}$, $E[z_{t-1}u_{t}]=\mathbf{0}_{K\times 1}$, $0<E[|x_{k,t-1}|]<\infty$ for all $k$, and that the matrix $E[z_{t-1}x_{t-1}^{\prime}]$ is invertible. [^10] Even when $\operatorname{sign}(x_{t-1})$ is constant, the univariate Cauchy estimator and associated tests remain well defined. In this case, the estimator simplifies to $\check{\beta}=\sum_{t=1}^{T}y_{t}\big/\sum_{t=1}^{T}x_{t-1}$, implying $\check{\beta}-\beta=\sum_{t=1}^{T}u_{t}\big/\sum_{t=1}^{T}x_{t-1}$, and Proposition 3.4 continues to hold. Hence, the lack of sign variation does not affect the validity of our methods in the univariate case. In the multiple-predictor case, however, the invertibility of $E[z_{t-1}x_{t-1}^{\prime}]$ imposes mild restrictions on sign changes across predictors. For practical applications, one can induce variation in the sign instrument by recentering the predictor, for example, $\tilde{x}_{t-1}=x_{t-1}-t^{-1}\sum_{s=1}^{t}x_{s-1}$, which preserves both the martingale structure and the asymptotic validity of the estimator. Then, the joint null hypothesis $H_{0}$ holds if and only if $H_{0}^{(k)}$ holds for all $k=1,\ldots,K$.
Lemma 4.1 justifies the use of multiple hypothesis testing based on univariate Cauchy estimators. [^11] See Harvey et al. (2015) for an application of the multiple-testing framework in predictive regressions, and Kostakis et al. (2015) for joint-predictability tests in the IVX framework. Note that the IVX approach may lose validity under heavy-tailed predictors or continuous-time data, whereas our method remains robust in such settings. In conjunction with the hybrid test introduced in Section 3.2, we compute the statistic $\tau(\check{\beta}_{k})$ for each parameter $\beta_{k}$, where $\check{\beta}_{k}$ denotes the corresponding Cauchy estimator. Let $p_{k}$ denote its $p$-value. The joint null hypothesis $H_{0}$ is rejected at level $\alpha$ if $\min_{k}p_{k}\leq\alpha/K$, following the Bonferroni correction.
This approach directly extends the univariate robust inference procedure to a multivariate setting and requires only mild conditions for the equivalence between $H_{0}$ and $\{H_{0}^{(k)}\}_{k=1}^{K}$. The Bonferroni correction imposes no assumptions on the joint distribution of the test statistics, which motivates its use here (see Holm, 1979; Benjamini and Hochberg, 1995; Shaffer, 1995).
We also note that under some additional regularity conditions, the joint hypothesis can be tested directly using a Wald-type statistic:
@@
and hence $W\to_{d}\chi_{K}^{2}$. [^12] As mentioned earlier, $z_{t}$ can be interpreted as an instrument. Therefore, one may use an alternative instrument, as in Shephard (2020), and construct a Wald-type test accordingly. We provide a detailed analysis of the Wald-type test in the Online Appendix, Section A.1.
While both the Wald test and the Bonferroni-adjusted marginal tests evaluate the same null hypothesis, they are not equivalent procedures except under special cases such as asymptotic orthogonality of the regressors. The Wald test exploits the full covariance structure and yields an elliptical rejection region, whereas the Bonferroni procedure yields a rectangular region and controls family-wise error. Consequently, the procedures may differ in finite-sample and asymptotic power when regressors are correlated. [^13] Again, our goal is not to design efficient procedures but to provide simple and robust methods that rely on minimal assumptions. We leave a systematic comparison between the Bonferroni-type multiple-testing procedure and the Wald-type joint test for future research.
@@ 4.2 Predictive Regressions with an Intercept
where $\bar{z}_{s}=s^{-1}\sum_{t=1}^{s}z_{t}$ for $z_{t}\in\{y_{t},x_{t-1},u_{t}\}$. However, $(u_{t}-\bar{u}_{T})$ is not a martingale difference sequence (MDS) with respect to $(\mathcal{F}_{t})$, invalidating the martingale CLT used in Sections 2 and 3. Specifically, the Cauchy estimator becomes
@@
and estimate this first-differenced (FD) model on alternating subsets of observations. [^15] Differencing $y_{t}$ to eliminate the nonzero mean is a widely used technique in the analysis of dynamic panel data with individual heterogeneity. See, among others, Anderson and Hsiao (1981); Anderson and Hsiao (1982); Arellano and Bond (1991). We focus on the even-indexed observations and define the modified Cauchy estimator:
@@
This estimator has two key properties. First, for even-indexed data, the regression error $u_{t}^{e}=u_{2t}-u_{2t-1}$ forms an MDS with respect to $\mathcal{F}_{t}^{e}:=\mathcal{F}_{2t}$ for $t=1,\ldots,T/2$. [^16] For odd-indexed data, $u_{t}^{o}=u_{2t+1}-u_{2t}$ forms an MDS with respect to $\mathcal{F}_{t}^{o}:=\mathcal{F}_{2t+1}$, yielding an analogous estimator $\check{\beta}_{o}$. Second, $\check{\beta}_{e}$ can again be viewed as an IV estimator, but it uses $\text{sign}(x_{t-2})$, which is $\mathcal{F}_{t-2}$-measurable, as the instrument. [^17] More generally, one may use $\text{sign}\!\left(\sum_{l\leq 2}w_{l}x_{t-l}\right)$ for deterministic weights $\{w_{l}\}$, provided $E[\text{sign}(\sum_{l\leq 2}w_{l}x_{t-l})(x_{t-1}-x_{t-2})]\neq 0$. Hence, $\text{sign}(x_{2t-2})(u_{2t}-u_{2t-1})$ is an MDS with respect to $(\mathcal{F}_{t}^{e})$.
@@
Although the odd-indexed estimator $\check{\beta}_{o}$ has analogous properties, $\check{\beta}_{e}$ and $\check{\beta}_{o}$ are typically dependent, with the dependence structure determined by the DGP of $(x_{t})$. Hence, unless additional assumptions are imposed, we restrict attention to a single subset of observations—either with even or odd indices. [^18] Using only half of the data is not uncommon in predictive regressions. See, for example, Zhu et al. (2014) and Liu et al. (2019), who employ long-lag differencing to eliminate intercepts. In addition, Dufour and Torrès (2000) uses a split-sample approach to address inference problems under a Markovian structure.
@@
$E[\operatorname{sign}(x_{t-1})x_{t}]\neq E[|x_{t-1}|].$
The condition may fail for certain unit-root processes. For instance, for a random walk $x_{t}=x_{t-1}+\varepsilon_{t}^{x}$, it does not hold. More generally, in the local-to-unity model of Phillips and Magdalinos (2007b),
@@
The first term diverges if and only if $\delta<1$ (see Lemma 3.2 of Phillips and Magdalinos, 2007b). Since $\varepsilon_{t}^{x}$ and $x_{t-1}$ may be dependent, it is generally the case that $E[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]\neq 0$, implying that the second term also diverges. Consequently, inconsistency arises only when $\delta=1$ and $E[\operatorname{sign}(x_{2t-2})\varepsilon_{2t-1}^{x}]=0$. In all other cases (i.e., $0\leq\delta<1$ or when the expectation is nonzero), the test remains consistent.
Moreover, the robustness to a nonzero intercept is not without cost. Specifically, the test based on $\tau(\check{\beta}_{e})$ is less powerful than that based on $\tau(\check{\beta})$ when the intercept is zero. A detailed power comparison is provided in Section A.2 of the Online Appendix.
@@ 5 Finite Sample Performance
This section investigates the finite-sample performance of the proposed methods. Two sets of simulation experiments are conducted. The first set is based on a continuous-time model and compares our robust $t$-statistic–based tests, $t_{q}(\check{\gamma})$ for $q\in\{8,12,16\}$, and the hybrid test $\tau(\check{\beta})$ with the Cauchy RT test of Choi et al. (2016) and the Cauchy VC test of Ibragimov et al. (2024). In this setting, the intercept is suppressed in the test statistics. The second set is based on a discrete-time predictive regression model and compares our procedures, $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, with the IVX test of Kostakis et al. (2015). In this setting, the statistics are computed with the intercept included in the estimation. [^19] The number of replications is 10,000 for each setting.
@@ 5.1.1 Simulation Design
Following Choi et al. (2016) and Ibragimov et al. (2024), we consider the continuous-time predictive regression model
@@
where $V_{t}$ and $W_{t}$ are Brownian motions with $E(V_{t}W_{t})=-0.98t$. The constant term in the predictive regression is set to zero, and recursive demeaning is applied as Choi et al. (2016). The model is observed at interval $\Delta=1/252$, corresponding to daily observations, so that a sample of length $T$ years contains $252T$ observations.
@@
  • CNST (Constant vvolatility): $\sigma_{t}^{2}=\sigma_{0}^{2}$, with $\sigma_{0}=1$.
@@
  • GBM (Geometric Brownian Motion): $d\sigma_{t}^{2}=\frac{1}{2}\frac{\bar{\omega}^{2}}{T}\sigma_{t}^{2}\,dt+\frac{\bar{\omega}^{2}}{\sqrt{T}}\sigma_{t}^{2}\,dZ_{t}$, where $Z_{t}$ is a Brownian motion correlated with $W_{t}$ such that $E(W_{t}Z_{t})=-0.4t$ and $\bar{\omega}=9$.
Following Choi et al. (2016), we set $T\in\{20,50,100\}$ (corresponding to 240, 600, and 1200 monthly observations) and $\bar{\kappa}\in\{0,5,10\}$ for the persistence parameter in (11). We consider a two-sided test of $H_{0}:\beta=0$ against $H_{A}:\beta\neq 0$.
@@ 5.1.2 Results
We first assess the empirical size of each test under $H_{0}:\beta=0$. The results for the four volatility models (CNST, SB, RS, and GBM) are reported in Table 1. Overall, both the $t$-statistic–based tests and the hybrid method exhibit satisfactory size performance, closely matching the nominal levels and performing comparably to the Cauchy RT and Cauchy VC tests. Among the $t$-based procedures, moderate partition numbers ($q=12$ or $16$) provide the most stable results, whereas smaller $q$ values tend to be mildly undersized. In the GBM case, where volatility is endogenously persistent, the $t$-statistic–based tests become slightly conservative but remain competitive with the Cauchy RT and VC methods.
Next, we analyze the finite-sample power properties of the tests. We consider $\beta\in\{0.004k,k=1,\cdots,5\}$ under the same volatility specifications. The results are summarized in Tables 2S.8. The proposed tests exhibit power comparable to that of the Cauchy RT and Cauchy VC procedures. For small samples ($T=20$), the Cauchy RT and VC tests occasionally show higher power, but the difference diminishes as $T$ increases. In certain settings, our methods even outperform the existing approaches. For instance, $t_{16}(\check{\gamma})$ dominates under $\beta=0.02$, $\bar{\kappa}=0.5$, and regime switching volatility (Table 4), whereas the hybrid test $\tau(\check{\beta})$ performs best under $\beta=0.004$, $\bar{\kappa}=20$, $T=20$, and regime switching volatility.
@@ 5.2.1 Simulation Design
We now examine the finite-sample performance of the proposed tests in a discrete-time setting with an intercept, comparing them to the IVX test of Kostakis et al. (2015). The DGP is specified as
$y_{t}$ $=\beta x_{t-1}+\sigma_{t}\varepsilon_{t},$ (12)
$x_{t}$ $=\left(1-\frac{\bar{\kappa}}{T}\right)x_{t-1}+\sigma_{t}\eta_{t},$
for $t=2,\ldots,12T$, where $T\in\{20,50,100\}$ corresponds to 20, 50, and 100 years of monthly data (i.e., 240, 600, and 1200 monthly observations, respectively). The constant term in the predictive regression is set to zero. We set $\beta\in\{0.5k:k=0,1,\ldots,5\}$ and $\bar{\kappa}\in\{0,50,100\}$, and consider a one-sided test of $H_{0}\!:\beta=0$ against $H_{A}\!:\beta>0$. [^20] The IVX test of Kostakis et al. (2015) performs well in two-sided testing for a broad class of models with constant volatility. However, as shown in Demetrescu et al. (2023), the IVX method exhibits severe size distortions in one-sided tests when regressors are highly persistent and endogenous. For this reason, we focus on the one-sided case to highlight the performance of our methods in this setting. Simulations for two-sided and joint hypothesis tests are also reported in Section B of the Online Appendix. Notably, while the IVX approach demonstrates reliable size control under constant volatility, it exhibits significant size distortions in the presence of non-constant volatility. In contrast, our methods remain robust to such heteroskedasticity.
@@
$\eta_{t}=\sum_{j=0}^{q}C_{j}v_{t-j},$ (13)
where $(\varepsilon_{t},v_{t})$ are jointly normal with correlation 0.98-0.98. In our simulations, we consider an MA(1) process with $C_{0}=C_{1}=1/\sqrt{2}$. [^21] Additional simulation results with an MA(3) process are reported in the Online Appendix, Section B.1. Moreover, as discussed in Section 4.2, the intercept-robust version of our test, $\tau(\check{\beta}_{e})$, becomes inconsistent when $x_{t}$ follows the local-to-unity process (10) with $\delta=1$ and $\eta_{t}$ is serially independent. For this reason, we do not consider i.i.d. $\eta_{t}$ in our simulations. See the Online Appendix, Section A.2, for a detailed theoretical power analysis of the test statistic $\tau(\check{\beta}_{e})$. The volatility process $\sigma_{t}$ follows the same specifications as in the continuous-time simulations, except that the GBM model is excluded.
We implement the hybrid tests based on the even and odd observations, denoted by $\tau(\check{\beta}_{e})$ and $\tau(\check{\beta}_{o})$, respectively (see (9)). For comparison, we also include the IVX test of Kostakis et al. (2015), in which the intercept is included in the estimation.